OpenStudy (anonymous):

-2 3/8 expressed as an improper fraction?

Mehek (mehek14):

-2*8 + 3 = ?

OpenStudy (anonymous):

Is it -19/8?

Mehek (mehek14):

-16 + 3 = -13 13/8

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wait it doesn't have 13/8 as a choice

Mehek (mehek14):

-13/8

OpenStudy (anonymous):

nor does it have that as a choice

Mehek (mehek14):

what does it have o-o

OpenStudy (anonymous):

-6/16, -19/8, -3/4, and 19/8

Mehek (mehek14):

wait is it -2 -3/8?

OpenStudy (anonymous):

-2 3/8

Mehek (mehek14):

3 is positive?

OpenStudy (anonymous):

ye you must have read it wrong

Mehek (mehek14):

no it should be -13/8 O_O -2 * 8 = -16 -16 + 3 = -13

OpenStudy (anonymous):

hm then idk k12 must've done something wrong?

Mehek (mehek14):

ye .. @pooja195 check for me

pooja195 (pooja195):

Looks fine to me :/

Mehek (mehek14):

-13/8 right?

pooja195 (pooja195):

Yeah

hero (hero):

-(2+3/8)

OpenStudy (anonymous):

Well if it is right then k12 doesn't have it as an answer choice

pooja195 (pooja195):

Is -19/8 a choice?

OpenStudy (anonymous):

yes

pooja195 (pooja195):

That is the correct answer then i cant think of another way to get another answer Since its a negative fraction you would subtract the 3 i think so -8 x 2= -16-3=-19/8 <---denominator stays the same

OpenStudy (anonymous):

should I put -19/8?

pooja195 (pooja195):

I would say that that's right provided the my explanation about :P

OpenStudy (anonymous):

Ok thanks

pooja195 (pooja195):

yw